State True or False : If for a given voltage , resistance is changed from R to R/n, power consumed changes from P to nP
1
Expert's answer
2012-07-05T10:11:07-0400
P = V · I = V · V / R = V^2 / R
P0 = V^2 / R
P1 = V^2 / (R/n) = n * V^2 /
R

Answer: true, power consumed changes from P to nP.